Find the hypotenuse of a right angle it its two legs are 6 and 8?
Answer by stanbon(75887)

Find the hypotenuse of a right angle it its two legs are 6 and 8?
----
h^2 = 6^2+8^2
---
h^2 = 36+64
---
h^2 = 100
---
Positive solution:
h = 10
==============
Cheers,
Stan H.